OpenStudy (anonymous):

What is the equation of a line passing through (-8, -3) and having a slope of -4

OpenStudy (math2400):

do u know what equation u can use?

OpenStudy (math2400):

y-y=m(x-x)

OpenStudy (math2400):

simply plug in your values. y'=-3 and x'=-8 m=-4

OpenStudy (math2400):

so u have y-(-3)=-4(x-(-8)) then just clean up and u get: y+3=-4x-32 and put it into slope intercept form by subtracting your 3: y=-4x-35